# Serial and Parallel Two-Column Probing for Mixed-Integer Programming

### v1

@@ 1 Introduction
For mixed integer programming (MIP) problems, presolving (or preprocessing) is a set of techniques run prior to the main branch-and-bound search that aims to transform a given instance into an equivalent but (hopefully) easier-to-handle formulation. Techniques include eliminating redundancies and strengthening the formulation [6, 24]. Presolve plays an important role in practical results for MIP solvers [8], as well as linear programming (LP) [9, 29], and mixed integer nonlinear programming (MINLP) [42, 39]. Besides general solvers, there are also problem-specific preprocessing methods (see, e.g. [18, 16, 31]).
@@
Probing plays an important role in MIP presolves with demonstrated impact on solvers, e.g. [6]. This technique temporarily fixes binary variables to obtain a variety of useful implications such as variable bound strengthening, and aggregations [45, 7, 2, 48, 32, 26, 1, 6]. The MIP version of probing follows probing for SAT solvers [37], and has subsequently also found use in extensions such as MINLP [35] and multi-stage integer programming [38, 43].
@@
The standard approach to probing involves fixing one binary variable at a time and running e.g. domain propagation [6, Sec. 7.2] and [2, Sec. 10.6]. Probing all binary variables can be potentially time-consuming due to the quadratic runtime complexity associated with the number of nonzero constraint coefficients; thus, solvers typically set a workload-limit and attempt to prioritize variables with a probing order [2, 4, 3, 6, 27]. In this paper, we propose a new two-column probing approach, involving pairs of variable per iteration. As with standard probing, in two-column probing we rely on workload-limits and judicious selection of (pairs of) variables. Our approach is a natural generalization of one-column probing that follows a similar philosophy applied to other parts of MIP such as: multiple columns to reduce models [14, 6]; multiple-row reduction [5]; matrix sparsification [17, 29, 6]; and two-row/column bound strengthening [26].
@@
Additionally, we develop a parallelization scheme for two-column probing that enables substantially more pairs to be processed, but may generate weaker (local) per-probing results compared to the (global) serial scheme. Parallelization has become increasingly important component of MIP solvers [15, 12, 8, 33]; we note also a growing literature in massively parallel computing for MIP (e.g. [41, 23, 46, 34, 47, 40]). Notably, Gleixner, Gottwald, and Hoen [27] parallelized one-column probing by distributing each task of probing a variable to a different thread. In their scheme, global variable bounds, aggregations, and the implication graph must be updated after each iteration; subsequently, the updates involve substantial (time-consuming) communication. For two-column probing, we develop a different approach in Sec. 4 that partitions the set of binary variables and distributes probing work across processors with lower communication costs. We note also that this paper shares the philosophy of our preceding work on parallel conflict graph management [21]: to develop intensive versions of existing preprocessing methods for which the overhead can be mitigated by parallel computing with carefully managed communication costs.
@@ 2.1 Notation and Basic Definitions
[…]
@@
Note that if $\circ_{i}$ is $\geq$, we can rewrite the constraint as $-A_{i\cdot}x\leq-b_{i}$, and if $\circ_{i}$ is $=$, we can split the constraint into two constraints $A_{i\cdot}x\leq b_{i}$ and $-A_{i\cdot}x\leq-b_{i}$. So we only consider constraints with the form of $A_{i\cdot}x\leq b_{i}$ in the remainder of this section. Let
@@ 2.2 Simple Pre-Presolves
At the preprocessing stage, given a $\mbox{MIP}=(\mathcal{M},\mathcal{N},\mathcal{I},A,b,c,\circ,\ell,u)$, we perform some pre-presolves to detect set packing constraints and constructing $CM$ and $CT$ for subsequent probing, which is described in Alg 1 and Alg 2. This applies well-studied techniques (see e.g. Achterberg et al. [6]).
@@ 9             Set $\mathcal{S}_{nsp}:=\mathcal{S}_{nsp}\cup\{A_{i\cdot}x\leq b_{i}\}$;
All techniques mentioned in line 2 of Alg. 1 can be found in [6, Sec 3.1 and 3.2]. Empty constraints and singletons are discarded as they cannot be probed; moreover, the strengthening of variables bounds allows for detection of binary variables that are otherwise encoded as general integers. The complexity of Alg 1 is $O(NNZ)$, where $NNZ$ is the number of nonzero elements in $A$ [27].
@@ 15             Set $CT[k]:=CT[k]\cup\{j\}$;
return $CM$, $CT$.
Algorithm 2 Building $CM$ and $CT$
@@
Specifically, because current solvers [30, 2, 25, 20] have already implemented Alg. 1 and it is fast, we excluded its runtime in experiments (which is marginally small regardless). However, we include the runtime of Alg. 2 because different solvers may construct clique table (or conflict graph [10, 1]) with different data structures, which makes various building runtime and may impact subsequent probing runtime. And $CM$ may not be used in other presolve routines.
@@ 3.1 Selecting Variable Pairs
When considering the order of variables to probe, Achterberg [2] determines a probing order of binary variables based on three factors: the number of constraints containing a binary variable $x_{j}$, the number of arcs of $x_{j}$ in the implication graph, and the number of conflicts between $x_{j}$ and other binary variables. For our implementation we focus on constructing an independent probing subroutine, and so we do not generate an implication graph prior to running probing; in practice the implication graph is generated from other preprocessing procedures and refined during probing. Thus, we define probing order in a simpler way, estimating the number of conflicts of binary variable $x_{j}$ by conflict number defined in Equation (3) by $CT$ from Alg. 2. The number of constraints containing $x_{j}$ is $\|A[:,i]\|_{0}$, where the zero-norm represents the number of nonzero elements from one vector. Because pairs of variables are considered, besides the number of constraints containing one binary variable, we consider the number of constraints containing two binary variables. This number is calculated via the coupling matrix $CM\in\mathbb{Z}_{+}^{|\mathcal{B}|\times|\mathcal{B}|}$ from Alg. 2.
Alg. 3 determines probing order. Line 1 selects Cand_Number candidate pairs of variables via binary search. $conflict$ in lines 3-8 acts a penalty to help avoid conflicting variables in the probing order—conflicts reduce the number of useful combinations among the pair of binary variables, i.e. $(0,0),(0,1),$ $(1,0),(1,1)$. Line 10 fetches up to a fixed number of pairs, Max_Probe_Number, for probing, and applies partial sorting (see e.g. [44]).
Input: $CM$, $CT$
@@
1 Set $Cand$ as an index set including indices of largest Cand_Number cells from $CM$;
@@ 6             Set $conflict=1$;
12Partially sort $Cand$ to extract the top Max_Probe_Number elements with $score(Cand_{i,j}^{(1)})\geq score(Cand_{i,j}^{(2)})\geq...\geq score(Cand_{i,j}^{(1000)})$;
@@ 3.2 Two Column Probing
Input: MIP, candidate pairs of variables $Cand$, $CT$, $ub,lb$
Output: $ub,lb$, $AF$, new conflicts set $New\_Conflict$
@@ 7             Call domain propagation to obtain implied bounds $ub(v),lb(v)$;
8             If the problem is infeasible, add $(2v_{1}-1)x_{i}+(2v_{2}-1)x_{j}\leq v_{1}+v_{2}-1$ to $CT$ and $New\_Conflict$;
@@ 15       Check termination condition;
Return $ub,lb$, $AF$, $New\_Conflict$.
@@
In the input, $Cand$ includes Max_Probe_Number pairs of binary variables selected from Alg. 3. Line 7 applies domain propagation (see e.g. [6, Sec. 3.2]) to get implied variable bounds. In line 8, $(2v_{1}-1)x_{i}+(2v_{2}-1)x_{j}\leq v_{1}+v_{2}-1$ cuts off $x_{i}=v_{1},x_{j}=v_{2}$, which is a detected clique. Line 10 fixes variables as follows:
@@
Aggregation and implication are then handled as in the standard one-variable probing [6, Sec. 7.2] or [2, Alg. 10.12, step 4 (c)(d)]. After line 12, we also add implications like $x_{i}=1,x_{j}=1\rightarrow x_{k}\geq b$ to $IG$. To avoid redundancy, pair implications are only added if the individual implications $x_{i}=1\rightarrow x_{k}\geq b$ or $x_{j}=1\rightarrow x_{k}\geq b$ are not in $IG$.
@@ 3.3 Discussion and Analysis
In terms of computational complexity, within the for-loop in line 2 of Alg. 4, domain propagation is in $O(NNZ)$, and updating variable bounds, $AF,IG,CT$ are in $O(n)$. Hence, the per-iteration complexity is the same as in one-column probing [27]. The maximum number of iterations is $O(|B|^{2})$ (governed by the number of possible pairs), whereas standard one-column probing is applied once per variable and so has $O(|B|)$ iterations; in either case, workload limits are imposed to make this a constant factor in practice. Note that $\lceil\frac{B}{2}\rceil$ pairs in $Cand$ are sufficient to attain the same (or better) results as one-column probing, provided such pairs do not share variables.
Practically speaking, the most time-consuming step in probing (both single- and two-column) is generating $CT,AF,IG$, and domain propagation. Alg. 4 calls $CT,AF,IG$, and domain propagation $4$ times, which is twice the computational effort of single-column probing, i.e. running single probing to each $x_{i}$ and $x_{j}$ separately. One benefit of this is that all aggregations and implications from single-column probing will also be detected in line 12. The reason is that the implied bounds in line 12 are no weaker than the implied bounds from the one-column probing. For the variable upper bound, suppose $v_{1},v_{2}\in\{0,1\}$,
@@
thus the implied upper bound from two-column-fixing is no weaker than the bound from one-column-fixing. The lower bound can be shown similarly. We also provide an example to show that two-column probing may return a stronger variable bound in domain propagation. Consider the following system:
$t+x+y+z\geq 1,$ (4)
$w+y\leq 1,$ (5)
$w-z\geq 0,$ (6)
$w,x,y,z\in\{0,1\},t\in[-3,3]$ (7)
Fixing $x=1$ implies $t\geq-2$ from Eq. (4). However, fixing either $x=1,w=1$ or $x=1,w=0$, implies $t\geq-1$, and so $x=1$ implies $t\geq-1$. Additionally, line 8 in Alg. 4 can detect some conflicts between binary variables and line 13 can provide two-variable implications that cannot be obtained by single-column probing.
However, although two-column probing can provide possibly stronger (and never weaker) reduction per-pair compared to one-column probing, not all variables are considered: in line 1 of Alg. 3, only pairs of variables appearing in at least one constraint together are considered. Thus, some binary variables that only appear in constraints alone are ignored by two-column probing; indeed, in experiments we observe that this constitutes nearly half of all binary variables. This is an implementation choice that allows us to easily test the effects of two-column probing above-and-beyond those of one-column probing.
@@ 3.3.1 Termination Conditions
[…]
@@ 3.4 Output Management
Two-column probing provides strengthened variable bounds, new conflicts, $AF$, and $IG$. We modify the original MIP by replacing the variable bounds with the strengthened bounds and adding clique cuts from new conflicts and aggregation constraints like $x_{i}=ax_{j}+b$ from $AF$. However, the number of implications from $IG$ is quite large, so we discard them; in a more mature implementation there may be a better management scheme that can pass $IG$ in whole or in part in a productive way.

### v2

@@ 1 Introduction
For mixed integer programming (MIP) problems, presolving (or preprocessing) is a set of techniques run prior to the main branch-and-bound search that aims to transform a given instance into an equivalent but (hopefully) easier-to-handle formulation. Techniques include eliminating redundancies and strengthening the formulation [achterberg2020presolve, PresolveProgress]. Presolve plays an important role in practical results for MIP solvers [achterberg2013mixed], as well as linear programming (LP) [PresolveLP, PresolveInteriorPoint], and mixed integer nonlinear programming (MINLP) [PresolveMINLP, PresolveMISDP]. Besides general solvers, there are also problem-specific preprocessing methods (see, e.g. [application1, application2, application3]).
@@
Probing plays an important role in MIP presolves with demonstrated impact on solvers, e.g. [achterberg2020presolve]. This technique temporarily fixes binary variables to obtain a variety of useful implications such as variable bound strengthening, and aggregations [FirstProbing, Achterberg2005, scip, aggregation, hoffman1993solving, TwoRowColumn, achterberg2007conflict, achterberg2020presolve]. The MIP version of probing follows probing for SAT solvers [lynce2003probing], and has subsequently also found use in extensions such as MINLP [ProbMINLP] and multi-stage integer programming [ProbStochastic, ProbStochastic2].
@@
The standard approach to probing involves fixing one binary variable at a time and running e.g. domain propagation [achterberg2020presolve, Sec. 7.2] and [scip, Sec. 10.6]. Probing all binary variables can be potentially time-consuming due to the quadratic runtime complexity associated with the number of nonzero constraint coefficients; thus, solvers typically set a workload-limit and attempt to prioritize variables with a probing order [scip, scip1, scip2, achterberg2020presolve, gleixner2023papilo]. In this paper, we propose a new two-column probing approach, involving pairs of variable per iteration. As with standard probing, in two-column probing we rely on workload-limits and judicious selection of (pairs of) variables. Our approach is a natural generalization of one-column probing that follows a similar philosophy applied to other parts of MIP such as: multiple columns to reduce models [ParallelRowColumn, achterberg2020presolve]; multiple-row reduction [MultipleRow]; matrix sparsification [MatrixSparse1, PresolveInteriorPoint, achterberg2020presolve]; and two-row/column bound strengthening [TwoRowColumn]. We highlight that two-variable probing was considered in Atamturk et al. [atamturk2000conflict] (see also [atamturk2000mixed]), deployed in the context of conflict graph construction—pairs of variables used to detect infeasible and also suboptimal assignments. In this paper we apply probing in a modern, integrated fashion: not only to detect conflicts but also can be propagated repeatedly for aggregations, bound strengthening and implications. Indeed, the clique table from a conflict graph procedure can be used to warmstart our probing orders.
@@
Additionally, we develop a parallelization scheme for two-column probing that enables substantially more pairs to be processed, but may generate weaker (local) per-probing results compared to the (global) serial scheme. Parallelization has become increasingly important component of MIP solvers [bixby2007progress, berthold2012solving, achterberg2013mixed, koch2022progress]; we note also a growing literature in massively parallel computing for MIP (e.g. [phillips2006massively, eckstein2015pebbl, shinano2016solving, koch2012could, shinano2018fiberscip, perumalla2021design]). Notably, Gleixner, Gottwald, and Hoen [gleixner2023papilo] parallelized one-column probing by distributing each task of probing a variable to different threads with a similar scheme. We note also that this paper shares the philosophy of our preceding work on parallel conflict graph management [parallelclqmerge]: to develop intensive versions of existing preprocessing methods for which the overhead can be mitigated by parallel computing with carefully managed communication costs.
@@ 2.1 Notation and Basic Definitions
with parameters $A\in\mathbb{R}^{m\times n}$, $c\in\mathbb{R}^{n}$, $b\in\mathbb{R}^{m}$, $\ell\in(\mathbb{R}\cup{-\infty})^{n}$, and $u\in(\mathbb{R}\cup{\infty})^{n}$. Variables $x\in\mathbb{R}^{n}$, and $x_{j}\in\mathbb{Z}$ for $j\in\mathcal{I}\subseteq\mathcal{N}=\{1,...,n\}$. Constraints with relations $\circ_{i}\in\{=,\leq,\geq\}$ for each row $i\in\mathcal{M}=\{1,...,m\}$. Furthermore, let $\mathcal{B}:=\{j\in\mathcal{I}\ |\ 0\leq x_{j}\leq 1\}$ be the set of indices for all binary variables. Let $NNZ$ denote the number of nonzero coefficients in $A$.
@@
Note that if $\circ_{i}$ is $\geq$, we can rewrite the constraint as $A_{i\cdot}x\leq-b_{i}$, and if $\circ_{i}$ is ==, we can split the constraint into two constraints $A_{i\cdot}x\leq b_{i}$ and $A_{i\cdot}x\leq-b_{i}$. So we only consider constraints with the form of $A_{i\cdot}x\leq b_{i}$ in the remainder of this section. Let
@@ 2.2 Simple Pre-Presolves
At the preprocessing stage, given a $\mbox{MIP}=(\mathcal{M},\mathcal{N},\mathcal{I},A,b,c,\circ,\ell,u)$, we perform some pre-presolves to detect set packing constraints and constructing $CM$ and $CT$ for subsequent probing, which is described in Alg 1 and Alg 2. This applies well-studied techniques (see e.g. Achterberg et al. [achterberg2020presolve]).
@@ 9      Set $\mathcal{S}_{nsp}:=\mathcal{S}_{nsp}\cup\{A_{i\cdot}x\leq b_{i}\}$;
All techniques mentioned in line 2 of Alg. 1 can be found in [achterberg2020presolve, Sec 3.1 and 3.2]. Empty constraints and singletons are discarded as they cannot be probed; moreover, the strengthening of variables bounds allows for detection of binary variables that are otherwise encoded as general integers. The complexity of Alg 1 is $O(NNZ)$, where $NNZ$ is the number of nonzero elements in $A$ [gleixner2023papilo].
@@ 15      Set $CT[k]:=CT[k]\cup\{j\}$;
return CMCM, CTCT.
Algorithm 2 Building CMCM and CTCT
@@
Specifically, because current solvers [gurobi, scip, copt, cplex2009v12] have already implemented Alg. 1 and it is fast, we excluded its runtime in experiments (which is marginally small regardless). However, we include the runtime of Alg. 2 because different solvers may construct clique table (or conflict graph [atamturk2000conflict, achterberg2007conflict]) with different data structures, which makes various building runtime and may impact subsequent probing runtime. And $CM$ may not be used in other presolve routines.
@@ 3.1 Selecting Variable Pairs
When considering the order of variables to probe, Achterberg [scip] determines a probing order of binary variables based on three factors: the number of constraints containing a binary variable $x_{j}$, the number of arcs of $x_{j}$ in the implication graph, and the number of conflicts between $x_{j}$ and other binary variables. For our implementation we focus on constructing an independent probing subroutine, and so we do not generate an implication graph prior to running probing; in practice the implication graph is generated from other preprocessing procedures and refined during probing. Thus, we define probing order in a simpler way, estimating the number of conflicts of binary variable $x_{j}$ by conflict number defined in Equation (3) by $CT$ from Alg. 2. The number of constraints containing $x_{j}$ is $\|A[:,i]\|_{0}$, where the zero-norm represents the number of nonzero elements from one vector. Because pairs of variables are considered, besides the number of constraints containing one binary variable, we consider the number of constraints containing two binary variables. This number is calculated via the coupling matrix $CM\in\mathbb{Z}_{+}^{|\mathcal{B}|\times|\mathcal{B}|}$ from Alg. 2.
Alg. 3 determines probing order. Line 1 selects Cand_Number candidate pairs of variables via binary search. conflictconflict in lines 3-8 acts a penalty to help avoid conflicting variables in the probing order—conflicts reduce the number of useful combinations among the pair of binary variables, i.e. $(0,0),(0,1),$ $(1,0),(1,1)$. Line 10 fetches up to a fixed number of pairs, Max_Probe_Number, for probing, and applies partial sorting (see e.g. [partialsort]).
Input: CMCM, CTCT
@@
1 Set CandCand as an index set including indices of largest Cand_Number cells from CMCM;
@@ 6      Set $conflict=1$;
12Partially sort CandCand to extract the top Max_Probe_Number elements with $score(Cand_{i,j}^{(1)})\geq score(Cand_{i,j}^{(2)})\geq...\geq score(Cand_{i,j}^{(1000)})$;
@@ 3.2 Two Column Probing
Input: MIP, candidate pairs of variables CandCand, CTCT, $ub,lb$
Output: $ub,lb$, AFAF, new conflicts set New_ConflictNew\_Conflict
@@ 7      Call domain propagation to obtain implied bounds $ub(v),lb(v)$;
8      If the problem is infeasible, add $(2v_{1}-1)x_{i}+(2v_{2}-1)x_{j}\leq v_{1}+v_{2}-1$ to CTCT and New_ConflictNew\_Conflict;
@@ 15   Check termination condition;
Return $ub,lb$, AFAF, New_ConflictNew\_Conflict.
@@
In the input, $Cand$ includes Max_Probe_Number pairs of binary variables selected from Alg. 3. Line 7 applies domain propagation (see e.g. [achterberg2020presolve, Sec. 3.2]) to get implied variable bounds. In line 8, $(2v_{1}-1)x_{i}+(2v_{2}-1)x_{j}\leq v_{1}+v_{2}-1$ cuts off $x_{i}=v_{1},x_{j}=v_{2}$, which is a detected clique. Line 10 fixes variables as follows:
@@
Aggregation and implication are then handled as in the standard one-variable probing [achterberg2020presolve, Sec. 7.2] or [scip, Alg. 10.12, step 4 (c)(d)]. After line 12, we also add implications like $x_{i}=1,x_{j}=1\rightarrow x_{k}\geq b$ to $IG$. To avoid redundancy, pair implications are only added if the individual implications $x_{i}=1\rightarrow x_{k}\geq b$ or $x_{j}=1\rightarrow x_{k}\geq b$ are not in $IG$.
@@ 3.3 Discussion and Analysis
3.3.1 Comparison with One-Column Probing
Iteration Complexity. In terms of computational complexity, within the for-loop in line 2 of Alg. 4, domain propagation is in $O(NNZ)$, and updating variable bounds, $AF,IG,CT$ are in $O(n)$. Hence, the per-iteration complexity is the same as in one-column probing [gleixner2023papilo]. The maximum number of iterations is $O(|B|^{2})$ (governed by the number of possible pairs), whereas standard one-column probing is applied once per variable and so has $O(|B|)$ iterations; in either case, workload limits are imposed to make this a constant factor in practice. Note that $\lceil\frac{B}{2}\rceil$ pairs in $Cand$ are sufficient to attain the same (or better) results as one-column probing, provided such pairs do not share variables.
Practically speaking, the most time-consuming step in probing (both single- and two-column) is generating $CT,AF,IG$, and domain propagation. Alg. 4 calls $CT,AF,IG$, and domain propagation $4$ times, which is twice the computational effort of single-column probing, i.e. running single probing to each $x_{i}$ and $x_{j}$ separately.
Strength. Two-column probing for a pair of variables weakly dominates the one-column equivalent in the following sense. All aggregations and implications from single-column probing will also be detected in line 12. The reason is that the implied bounds in line 12 are no weaker than the implied bounds from the one-column probing. For the variable upper bound, suppose $v_{1},v_{2}\in\{0,1\}$,
@@ 3.3.1 Comparison with One-Column Probing
thus the implied upper bound from two-column-fixing is no weaker than the bound from one-column-fixing. Dominance in the lower bound can be shown similarly.
The example below shows that two-column probing may return a stronger variable bound in domain propagation.
Example Consider the following system:
$t+x_{1}+x_{2}+x_{3}+x_{4}\geq 4,$ (4)
$y-x_{1}\leq 0,$ (5)
$y+x_{2}\geq 1,$ (6)
$z-x_{3}\leq 0,$ (7)
$z+x_{4}\geq 1,$ (8)
$y,z\in\{0,1\},t\in[0,4],x\in[0,1]^{4}.$ (9)
One-column probing probes $y$ and $z$ separately, and the variable bound of $t$ can be strengthened to $[0,3]$. Two-column Probing probes $y$ and $z$ together, and the variable bound of $t$ is updated to $[0,2]$, which is tighter.
Additionally, line 8 in Alg. 4 can detect some conflicts between binary variables and line 13 can provide two-variable implications that cannot be obtained by single-column probing.
Balancing Reduction and Runtime. Although two-column probing can provide possibly stronger (and never weaker) reduction per-pair compared to one-column probing, for practical purposes not all pairs are probed (even in the one-column case only a subset of variables are probed): in line 1 of Alg. 3, only pairs of variables appearing in at least one constraint together are considered. Thus, some binary variables that only appear in constraints alone are ignored by our two-column probing implementation; indeed, in experiments we observe that this constitutes nearly half of all binary variables. Avoiding one-column probes is an implementation choice that allows us to easily test the effects of two-column probing separate from those of one-column probing. The central empirical question, then, is whether one can choose pairs of variables to probe that provide sufficiently strengthened results above-and-beyond the single column (or no probing) counterfactual that can make up for the added compute.
@@ 3.3.2 Termination Conditions
We set three criteria to terminate two-column probing, adapting the one-column procedure from SCIP [scip, Algorithm 10.12]. First, we set a maximum iteration limit Max_Probe_Number, i.e. a limit on the number of pairs probed per thread. Second, we set a global time limit of $30$ seconds. Third, we set a soft stopping rule using a parameter $\texttt{eff}=0$ to estimate the computational effort expended. At the beginning of each iteration, we set $\texttt{eff}=0.9\times\texttt{eff}$, and at the end of each iteration, $\texttt{eff}=\texttt{eff}+110$. If two variables are fixed in one iteration, then $\texttt{eff}=0$; if only one variable is fixed, set $\texttt{eff}=0.5\times\texttt{eff}$. If any new conflicts or aggregations are generated, set $\texttt{eff}=0.8\times\texttt{eff}$. If we update the $IG$, $\texttt{eff}=0.9\times\texttt{eff}$. If the effort score exceeds a threshold eff_threshold, e.g. $\texttt{eff}>\texttt{eff\_threshold}$, then we stop the algorithm. These parameter values were taken from the analogous dynamic limits in SCIP’s one-column probing procedure and were not derived from empirical tuning.
@@ 3.4 Output Management
Two-column probing provides strengthened variable bounds, new conflicts, $AF$, and $IG$. We modify the original MIP by replacing the variable bounds with the strengthened bounds and adding clique cuts from new conflicts and aggregation constraints like $x_{i}=ax_{j}+b$ from $AF$. However, the number of implications from $IG$ is quite large, so we discard them; in a more mature implementation, there may be a better management scheme that can pass $IG$ in whole or in part in a productive way.